OpenStudy (anonymous):

Using complete sentences, explain which method you would use to solve the following system of equations and why. In your answer, include the solution to one of the variables and how you found it using the method you chose. x - 5y + 2z = 0 x + 4y - z = 12 2x - y + 3z = 10

OpenStudy (anonymous):

I would choose elimination just because there is 3 variables to find. x - 5y + 2z = 0 --->(-1) x + 4y - z = 12 ------------ -x + 5y - 2z = 0 (result of multiplying by -1) x + 4y - z = 12 ------------add 9y - 3z = 12 x + 4y - z = 12 -->(-2) 2x - y + 3z = 10 -------------- -2x - 8y + 2z = -24 (result of multiplying by -2) 2x - y + 3z = 10 -------------add -9y + 5z = -14 9y - 3z = 12 -9y + 5z = -14 ------------add 2z = -2 z = -1

OpenStudy (anonymous):

Thank u thank u thank u sooooooooooooooo much

OpenStudy (anonymous):

matrices would probably be easier, but I am not good with matrices. Glad I could help